OpenStudy (anonymous):

A wire loop has an area of 0.12 m^2, and the magnetic field through it increases from 0 to 0.20 T in 0.30 s. What is the induced emf in the loop? A. 0.80 V B. 0.080 V C. 0.024 V D. 0.0072 V

OpenStudy (michele_laino):

the magnetic flux change is: \[\large \Delta \Phi = \left( {0.2 - 0} \right) \times 0.12 = ...Weber\]

OpenStudy (anonymous):

0.024? choice C is our solution?

OpenStudy (michele_laino):

no, since the magnitude of the requested emf is given by the subsequent formula: \[\large E = \frac{{\Delta \Phi }}{{\Delta t}} = \frac{{0.024}}{{0.3}} = ...volts\]

OpenStudy (anonymous):

ohh okay! so now we get 0.08?so it should be choice B is the solution?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

yay! thank you!!:)

OpenStudy (michele_laino):

:)